# Achieving millisecond coherence fluxonium through overlap Josephson junctions

### v1

@@ Abstract
Fluxonium qubits are recognized for their high coherence times and high operation fidelities, attributed to their unique design incorporating over 100 Josephson junctions per superconducting loop. However, this complexity poses significant fabrication challenges, particularly in achieving high yield and junction uniformity with traditional methods. Here, we introduce an overlap process for Josephson junction fabrication that achieves nearly 100% yield and maintains uniformity across a 2-inch wafer with less than 5% variation for the phase slip junction and less than 2% for the junction array. Our compact junction array design facilitates fluxonium qubits with energy relaxation times exceeding 1 millisecond at the flux frustration point, demonstrating consistency with state-of-the-art dielectric loss tangents and flux noise across multiple devices. This work suggests the scalability of high coherence fluxonium processors using CMOS-compatible processes, marking a significant step towards practical quantum computing.
@@ I Introduction
Fluxonium qubits [1] have garnered increasing interest owing to their significantly longer coherence times [2, 3, 4] and enhanced fidelity in gate operations [5, 6, 7, 8], positioning them as a viable alternative for the development of fault-tolerant superconducting quantum processors. This superior performance is attributed to the intrinsic characteristics of their energy spectrum. In particular, the fluxonium spectrum exhibits a first-order insensitivity to external flux variations at the flux frustration point, effectively creating a dephasing-resistant sweet spot. At this juncture, the qubit’s energy gap $f_{01}$ is determined by the quantum tunneling amplitude between two fluxon states, which are delineated within a double-well potential. This quantum tunneling amplitude is exponentially suppressed by the Josephson energy barrier, relegating fluxonium qubits to operate at significantly lower frequencies ($f_{01}<1$ GHz) in comparison to transmon qubits [9]. Consequently, this operational frequency reduction bolsters the qubits’ $T_{1}$ lifetime, assuming a similar quality factor $Q=2\pi f_{01}T_{1}$. Critically, the diminution in qubit frequency does not translate to a compromise in operational speed. The inherent substantial anharmonicity of the fluxonium spectrum facilitates the execution of rapid and high-fidelity operations either within the computational subspace [5, 6, 7, 10] or by transiently or virtually engaging the noncomputational states [11, 8, 12].
@@
The advantageous characteristics of fluxonium qubits are offset by their increased fabrication complexity relative to transmon qubits. The simplicity of transmon qubits allows for the widespread adoption of double-angle evaporation techniques to construct Josephson junctions (JJ) with relatively reliable process control and accurate parameter targeting [13, 14, 15, 16, 17]. In contrast, the fabrication of each fluxonium qubit not only involves creating a phase slip JJ but also incorporating an array of approximately 100 JJs to function as a large shunting inductor. This complexity introduces substantial challenges for device yield through increased variability in qubit parameters. Two primary issues exacerbate parameter variability in fluxonium qubit fabrication: the impact of film surface and pattern edge roughness, as well as sidewall deposition effects due to angle evaporation, both of which constrain the accuracy of phase slip JJ parameters similar to transmons. Additionally, the effective junction area in Josephson junction arrays (JJAs) is highly sensitive to the evaporation angle, which varies significantly across the wafer surface, complicating consistent fabrication on larger wafers.
@@
To address this limitation and facilitate scalable fluxonium processor production, we developed an overlap junction process for fluxonium qubit fabrication [18, 19, 20]. In the overlap junction process flow, no wafer tilting is required for junction electrodes deposition, and the top and bottom electrode formations are performed in two separated patterning and deposition steps, with a vacuum break in between. This overlap approach has exhibited nearly 100% yield and uniformity across a 2-inch wafer with less than 5% variation for the phase slip junction and less than 2% for the junction array. Across multiple fluxonium devices, we reached dielectric loss tangent $\tan\delta_{C}\in(1.2,5.0)\times 10^{-6}$ and $1/f$ flux noise amplitude $A_{\Phi}\in(1.4,2.6)~{}\mu\Phi_{0}/\sqrt{\text{Hz}}$; here $\Phi_{0}=h/(2e)$ is the flux quantum. On our best device, we achieved $T_{1}=1.168$ ms and $T_{2,\mathrm{echo}}=0.943$ ms at the qubit’s flux frustration position with the qubit frequency $f_{01}=197$ MHz. Compared with our previous work [21], we attribute the improvement on coherence times to the flux noise reduction due to the robust compact JJA enabled by our overlap process. The dielectric loss and flux noise levels are comparable with the best performing fluxonium qubits made with the angle-evaporation techniques and embedded in a 3-dimensional cavity, tailored for high coherence demonstration [2, 4]. Our work suggests the feasibility of scaling high coherence fluxonium processor with a CMOS-compatible fabrication process.
@@ II Fluxonium Qubits with overlap junctions
We choose 2-inch sapphire wafer as the substrate for our fluxonium processor due to its low dielectric loss [22]. Tantalum film is deposited on the sapphire substrate to form superconducting circuit elements except for the formation of the qubit loop. As shown in Fig. 1a-d, our fluxonium qubit consists of a single phase slip Josephson junction (referred as JJ hereafter) connected in parallel with a shunting capacitor and an array of 100 Josephson junctions as the shunting inductor. In this work, we adopt an overlap junction process combined with lift-off technique (see Supplementary Material for detail) to fabricate JJ/JJA with Al/AlOx/Al structure. The process starts with an Al deposition to form the bottom electrode (BE), followed by e-beam resist lift-off and top electrode (TE) patterning. The wafer is then loaded into vacuum chamber for in situ barrier AlOx formation and TE Al deposition. Prior to AlOx formation, we implement a well calibrated argon (Ar) milling process to remove native oxides for AlOx barrier quality control.
@@
To evaluate qubit yield and process variability, we measure normal resistance ($R_{n}$) of JJ and JJA with various junction sizes on 2-inch test structure wafers. Each test wafer contains 16 dies with 16 JJ or JJA of various sizes on each die. Excluding a few data points from wafer-edge dies with known pattern issues, almost all of the JJ/JJA structures show expected $R_{n}$ values from wafers we measured, which is an indication of near perfect qubit yield on our device wafers. For the phase slip JJs, the across wafer relative standard deviation (RSD%) of $R_{n}$ trends down with increasing junction size, from 8.5% at 0.01 $\mu m^{2}$ to 2.8% at 0.42 $\mu m^{2}$. For JJA with single junction size of 0.5 $\mu m^{2}$ to 1 $\mu m^{2}$, RSDis lower at 1.5-2.5% compared to smaller phase slip JJ, shown in Fig. 2d. In our fluxonium design, the typical junction dimensions for JJ and JJA are 0.05 $\mu m^{2}$ and 1 $\mu m^{2}$, with RSDbeing 5% and 1.5% respectively. Transmission electron microscope (TEM) cross section of a typical single phase slip JJ made with the overlap process is shown in Fig. 2a. The thickness target of BE Al film is 60 nm, and it came out at 45 nm in the final structure, which means the Ar ion milling consumes 15 nm of Al prior to barrier AlOx formation. The compositional analysis performed with the energy dispersive X-ray spectroscopy (EDX) of the TEM shows mostly Al concentration in the junction with O occurring in a smooth interface between top and bottom electrodes, as depicted in Fig. 2b and c.
@@
It is also worth mentioning that we have been implementing the overlap junction process in our 22-qubit fluxonium processor fabrication on 4-inch sapphire wafers (see Supplementary Material for the statistics of the test structures). At the time of writing, two 22-qubit fluxonium chips from different wafers have been fully measured and characterized, and all qubits show normal operation with relatively accurate qubit parameters. This is strong evidence that the overlap junction process can be easily transferred to even larger wafer scale without any obvious issue. However, the process flow presented in this study, which utilizes e-beam lithography and lift-off techniques for JJ/JJA formation, is not immediately compatible with the contemporary 300 mm semiconductor manufacturing process. Future research will investigate the use of photolithography and subtractive etching for JJ/JJA formation, aiming to achieve a CMOS-compatible process flow for fluxonium processors.
@@ III Demonstration of high coherence fluxonium
In Fig. 3a, we present the measured energy relaxation ($T_{1}$) and dephasing ($T_{2}$) processes of one qubit at its flux frustration position. In the $T_{1}$ measurement, we initialize the qubit in either the ground |$\ket{0}$ or excited |$\ket{1}$ state at the beginning of each experiment via projective readout and herald the desired initial state (see Supplementary Material). The postselected population decay from |$\ket{1}$ and |$\ket{0}$ is jointly fitted with two exponential decays of the same decay constant $T_{1}$ and thermal state population $b$, i.e. $(P_{1,\ket{1}},P_{1,\ket{0}})=(a_{1}\exp(-t/T_{1})+b,a_{2}\exp(-t/T_{1})+b)$, where $P_{i,\ket{j}}$ means measuring the |$\ket{i}$ population with qubit initialized at |$\ket{j}$. We obtain $T_{1}=1.168$ ms and estimate the qubit’s effective temperature $T=18.7$ mK from the thermal state population. In the $T_{2}$ measurement, we utilize the spin echo pulse sequence after qubit initialization. For each delay, we measure the magnitude of the Bloch vector projected in three equally spaced phase angles to extract a characteristic dephasing time $T_{2,\mathrm{echo}}=0.943$ ms, from an exponential fit. Assuming photon shot noise as the dominant source of dephasing, we derive an average of $4\times 10^{-3}$ residue photons in the readout cavity (see Supplementary Material).
@@ IV Dielectric loss and flux noise
Figure 4: $\bf{a}$ Relaxation times $T_{1}$ versus $\Phi_{\mathrm{ext}}$ for qubit G, with a minimum frequency ($f_{01}$) of 197 MHz. Solid and dashed lines represent the dielectric loss model (Eq. 1) predictions at $T=0$ and $T=18.7$ mK, respectively. Inset: Qubit spectrum near flux frustration point. $\bf{b}$ Dephasing times $T_{2}$ versus $\Phi_{\mathrm{ext}}$ for the same qubit, using free induction (Ramsey) and various dynamical decoupling sequences (spin echo, CPMG(2), CPMG(4), CPMG(8)). The black lines indicate dephasing times due to flux noise derived from a joint fit of all data. $\bf{c}$ $\zeta$ versus $f_{01}$ for our qubit (blue dots) and the $E_{J}$-tunable qubit from Ref. [21] (orange dots). Green and red solid lines represent $\zeta(f_{01})$ calculated using the dielectric loss model (Eq. 1) at $T=18.7$ mK and $T=0$, respectively. Predictions using the flux noise model different $A_{\Phi}$ are shown as dashed lines. The grey shaded area signifies predicted $\zeta(f_{01})$ combining dielectric loss ($\tan\delta_{C}\in(1.21,4.93)\times 10^{-6}$ at $T=0$) and flux noise at $A_{\Phi}=2.43~{}\mu\Phi_{0}/\sqrt{\text{Hz}}$.
@@
In Fig. 4a and b, we present the measured $T_{1}$ and $T_{2}$ times of qubit G as functions of $\Phi_{\text{ext}}$, respectively. Each data point for $T_{1}$ is extracted from fitting the relaxation data with an exponential decay curve at a specific flux position. The $T_{1}$ versus $\Phi_{\text{ext}}$ data exhibit a plateau near the flux frustration position and demonstrate suppression at $\Phi_{\text{ext}}=\Phi_{0}/2\pm 0.005\Phi_{0}$, likely due to interaction with a TLS at the corresponding qubit frequency. This flux dependence of $T_{1}$ precludes inductive-like loss channels [26]. This data is subsequently fitted using two variants of the dielectric loss model to determine the dielectric loss tangent, $\tan\delta_{C}$. One variant adheres to the widely used phenomenological dielectric loss model [2, 3]:
@@
dependent on the qubit’s finite effective temperature $T$. The alternative model posits a bath of TLS weakly interacting with the qubit, demonstrating temperature independent loss by taking the saturation effect of the TLS at finite temperature into account [21]. This model coincides with Eq. 1 at $T\rightarrow 0$, where the $\coth$ term converges to unity. The choice of model influences the derived loss tangent $\tan\delta_{C}$, typically yielding a higher value for the TLS bath model when the qubit energy $\hbar\omega_{01}\lesssim k_{B}T$. The qubit parameters such as the matrix element $\langle 0|\hat{\varphi}|1\rangle$, the charging energy $E_{C}$, and the qubit frequency $\omega_{01}=2\pi f_{01}$ are independently obtained from the qubit spectroscopy, where a representative spectrum of the corresponding qubit around the flux frustration position ($\Phi_{\text{ext}}=\Phi_{0}/2$) are shown in the inset of the Fig. 4a. We observe occasional suppression of $T_{1}$ at $\Phi_{\text{ext}}=\Phi_{0}/2$, suggestive of a TLS drifting into resonance with the qubit, when the qubit is idled at the same specific external flux position ($\Phi_{\text{idle}}=\Phi_{0}/2$). This suppression, however, can be mitigated by idling the qubit at $\Phi_{\text{idle}}\neq\Phi_{0}/2$ (see Supplementary Material for details). For instance, the data in Fig. 4a is collected with the qubit idled at $\Phi_{\text{idle}}=0.505\Phi_{0}$. This behavior might stem from alterations in the TLS environment triggered by qubit dynamics, presenting an intriguing avenue for future research.
In Fig. 4b, we present the typical qubit dephasing times $T_{2}$ as a function of $\Phi_{\text{ext}}$, obtained from different dynamical decoupling schemes that utilize Ramsey, spin echo, CPMG($N$[27, 28] pulse sequences (with $N=2,4,8$ representing the number of $\pi$ pulses). Each data point is the result of fitting a Gaussian decay to the dephasing [29, 30, 31]. At the flux frustration point ($\Phi_{\text{ext}}=\Phi_{0}/2$), where the qubit exhibits first-order insensitivity to flux noise, we record the maximum $T_{2}$. This dephasing time $T_{2}$ decreases monotonically with deviation from this optimal position, a behavior typical of flux-type qubits [29]. The complete $T_{2}$ versus $\Phi_{\mathrm{ext}}$ dataset, including Ramsey, spin echo, and CPMG($N$) sequences, is then fitted to a $1/f$ flux noise model [30], i.e. $S_{\Phi}(\omega)=A_{\Phi}^{2}(2\pi\times 1~{}\text{Hz}/\omega)$ for the extraction of the flux noise amplitude $A_{\Phi}$ (see Supplementary Material for more details).
@@
Table 1: Summary of device parameters and coherence metrics. $T_{1}$ and $T_{2,\mathrm{echo}}$ values, are obtained by fitting relaxation and dephasing to exponential decays at the flux frustration position. Asterisks (*) highlight observations of non-exponential $T_{1}$ decay. Coherence times versus $\Phi_{\text{ext}}$ for qubit H were not collected.

### v2

# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

@@ Abstract
Fluxonium qubits are recognized for their high coherence times and high operation fidelities, attributed to their unique design incorporating a superinductor, which is typically implemented using an array of over 100 Josephson junctions. However, this complexity poses significant fabrication challenges, particularly in achieving high yield and junction uniformity with traditional methods. Here, we introduce an overlap process for Josephson junction fabrication that achieves nearly 100% yield and maintains uniformity across a 2-inch wafer with less than 5% variation for the phase slip junction and less than 2% for the entire junction array. We use a compact junction array design that achieves state-of-the-art dielectric loss tangents and flux noise levels, as confirmed by multiple devices. This enables fluxonium qubits to reach energy relaxation times exceeding 1 millisecond at the flux frustration point. This work paves the way for scalable high coherence fluxonium processors using CMOS-compatible processes, marking a significant step towards practical quantum computing.
@@ I Introduction
Fluxonium qubits [1] have garnered increasing interest owing to their significantly longer coherence times [2, 3, 4] and enhanced fidelity in gate operations [5, 6, 7, 8], positioning them as a viable alternative for the development of fault-tolerant superconducting quantum processors. This superior performance is attributed to the intrinsic characteristics of their energy spectrum. In particular, the fluxonium spectrum exhibits a first-order insensitivity to external flux variations at the flux frustration point, effectively creating a dephasing-resistant sweet spot. At this juncture, the qubit’s energy gap $f_{01}$ is determined by the quantum tunneling between two fluxon states within a double-well potential. This tunneling, known as coherent quantum phase slip [9], has its amplitude exponentially suppressed by the Josephson energy barrier of the small phase slip junction. As a result, fluxonium qubits operate at significantly lower frequencies ($f_{01}<1$ GHz) compared to transmon qubits [10]. Consequently, this operational frequency reduction bolsters the qubits’ $T_{1}$ lifetime, assuming a similar quality factor $Q=2\pi f_{01}T_{1}$. Critically, the diminution in qubit frequency does not translate to a compromise in operational speed. The inherent substantial anharmonicity of the fluxonium spectrum has inspired a variety of proposals for executing high-fidelity operations [11, 12, 13, 14, 15, 16]. Building on these proposals, significant advancements have been made in implementing rapid and high-fidelity two-qubit gates [17, 5, 18, 7, 8, 19, 6], as well as in readout [20] and reset techniques [21, 22].
@@
The advantageous characteristics of fluxonium qubits are offset by their increased fabrication complexity relative to transmon qubits. The simplicity of transmon qubits allows for the widespread adoption of double-angle evaporation techniques to construct Josephson junctions (JJ) with relatively reliable process control and accurate parameter targeting [23, 24, 25, 26, 27]. In contrast, the fabrication of each fluxonium qubit not only involves creating a phase slip JJ (orange box in Fig. 1b) but also incorporating an array of approximately 100 JJs (green box in Fig. 1b) to function as a large shunting inductor. This complexity introduces substantial challenges for device yield through increased variability in qubit parameters. Two primary issues exacerbate parameter variability in fluxonium qubit fabrication: the impact of film surface and pattern edge roughness, as well as sidewall deposition effects due to angle evaporation, both of which constrain the accuracy of phase slip JJ parameters similar to transmons. Additionally, the effective junction area in Josephson junction arrays (JJAs) is highly sensitive to the evaporation angle, which varies significantly across the wafer surface, complicating consistent fabrication on larger wafers.
@@
To address this limitation and facilitate scalable fluxonium processor production, we developed an overlap junction process for fluxonium qubit fabrication [28, 29, 30]. In the overlap junction process flow, no wafer tilting is required for junction electrodes deposition, and the top and bottom electrode formations are performed in two separated patterning and deposition steps, with a vacuum break in between. This overlap approach has exhibited nearly 100% yield and uniformity across a 2-inch wafer with less than 5% variation for the phase slip junction and less than 2% for the junction array. The variation in the JJA is characterized by the variation in the total normal resistance $R_{n}$ of all the junctions in series within the entire array. Across multiple fluxonium devices, we reached dielectric loss tangent $\tan\delta_{C}\in(1.2,5.0)\times 10^{-6}$ and $1/f$ flux noise amplitude $A_{\Phi}\in(1.4,2.6)~{}\mu\Phi_{0}/\sqrt{\text{Hz}}$; here $\Phi_{0}=h/(2e)$ is the flux quantum. On our best device, we achieved $T_{1}=1.168$ ms and $T_{2,\mathrm{echo}}=0.943$ ms at the qubit’s flux frustration position with the qubit frequency $f_{01}=197$ MHz. Compared with our previous work [31], we attribute the improvement on coherence times to the flux noise reduction due to the robust compact JJA enabled by our overlap process. The dielectric loss and flux noise levels are comparable with the best performing fluxonium qubits made with the angle-evaporation techniques and embedded in a 3-dimensional cavity, tailored for high coherence demonstration [2, 4]. Our work suggests the feasibility of scaling high coherence fluxonium processor with a CMOS-compatible fabrication process.
@@ II Fluxonium Qubits with overlap junctions
We choose 2-inch sapphire wafer as the substrate for our fluxonium processor due to its low dielectric loss [32]. Tantalum film is deposited on the sapphire substrate to form superconducting circuit elements except for the formation of the qubit loop. As shown in Fig. 1a-d, our fluxonium qubit consists of a single phase slip Josephson junction (referred as JJ hereafter) connected in parallel with a shunting capacitor and an array of 100 Josephson junctions as the shunting inductor. In this work, we adopt an overlap junction process combined with lift-off technique (see Appendix A) to fabricate JJ/JJA with Al/AlOx/Al structure. The process starts with an Al deposition to form the bottom electrode (BE), followed by e-beam resist lift-off and top electrode (TE) patterning. The wafer is then loaded into vacuum chamber for in situ barrier AlOx formation and TE Al deposition. Prior to AlOx formation, we implement a well calibrated argon (Ar) milling process to remove native oxides for AlOx barrier quality control.
@@
To evaluate qubit yield and process variability, we measure normal resistance ($R_{n}$) of JJ and JJA with various junction sizes on 2-inch test structure wafers. The yield is defined as no hard failure (opens/shorts) of qubits on test patterns, since real devices/qubits cannot be measured without any damage. Each test wafer contains 16 dies with 16 JJ or JJA of various sizes on each die. Excluding a few data points from wafer-edge dies with photolithography issues, almost all of the JJ/JJA structures show expected $R_{n}$ values from wafers we measured, which is an indication of near perfect qubit yield on our device wafers. For the JJs, the across wafer relative standard deviation (RSD%) of $R_{n}$ trends down with increasing junction size, from 8.5% at 0.01 $\mu m^{2}$ to 2.8% at 0.42 $\mu m^{2}$. For JJA with single junction size of 0.5 $\mu m^{2}$ to 1 $\mu m^{2}$, RSDis lower at 1.5-2.5% compared to smaller phase slip JJ, shown in Fig. 2d. We emphasize that the RSDof the JJA is measured by the within-wafer variation in the normal resistance $R_{n}$ of the entire array, which characterizes the variation in the total inductance formed by the JJA. In our fluxonium design, the typical junction dimensions for JJ and JJA are 0.05 $\mu m^{2}$ and 1 $\mu m^{2}$, with RSDbeing 5% and 1.5% respectively. Transmission electron microscope (TEM) cross section of a typical single phase slip JJ made with the overlap process is shown in Fig. 2a. The thickness target of BE Al film is 60 nm, and it came out at 45 nm in the final structure, which means the Ar ion milling consumes 15 nm of Al prior to barrier AlOx formation. The compositional analysis performed with the energy dispersive X-ray spectroscopy (EDX) of the TEM shows mostly Al concentration in the junction with O occurring in a smooth interface between top and bottom electrodes, as depicted in Fig. 2b and c.
@@
It is also worth mentioning that we have been implementing the overlap junction process in our 22-qubit fluxonium processor fabrication on 4-inch sapphire wafers (see Appendix B for the statistics of the test structures). At the time of writing, two 22-qubit fluxonium chips from different wafers have been fully measured and characterized, and all qubits show normal operation with relatively accurate qubit parameters. This is strong evidence that the overlap junction process can be easily transferred to even larger wafer scale without any obvious issue. However, the process flow presented in this study, which utilizes e-beam lithography and lift-off techniques for JJ/JJA formation, is not immediately compatible with the contemporary 300 mm semiconductor manufacturing process. Future research will investigate the use of photolithography and subtractive etching for JJ/JJA formation, aiming to achieve a CMOS-compatible process flow for fluxonium processors.
@@ III Demonstration of high coherence fluxonium
In Fig. 3a, we present the measured energy relaxation ($T_{1}$) and dephasing ($T_{2}$) processes of one qubit at its flux frustration position. In the $T_{1}$ measurement, we initialize the qubit in either the ground |$\ket{0}$ or excited |$\ket{1}$ state at the beginning of each experiment via projective readout and herald the desired initial state (see Appendix D). The postselected population decay from |$\ket{1}$ and |$\ket{0}$ is jointly fitted with two exponential decays of the same decay constant $T_{1}$ and thermal state population $b$, i.e. $(P_{1,\ket{1}},P_{1,\ket{0}})=(a_{1}\exp(-t/T_{1})+b,a_{2}\exp(-t/T_{1})+b)$, where $P_{i,\ket{j}}$ means measuring the |$\ket{i}$ population with qubit initialized at |$\ket{j}$. We obtain $T_{1}=1.168$ ms and estimate the qubit’s effective temperature $T=18.7$ mK from the thermal state population. In the $T_{2}$ measurement, we utilize the spin echo pulse sequence after qubit initialization. For each delay, we measure the magnitude of the Bloch vector projected in three equally spaced phase angles to extract a characteristic dephasing time $T_{2,\mathrm{echo}}=0.943$ ms, from an exponential fit. Assuming photon shot noise as the dominant source of dephasing, we derive an average of $4\times 10^{-3}$ residue photons in the readout cavity (see Appendix G).
@@ IV Dielectric loss and flux noise
Figure 4: $\bf{a}$ Relaxation times $T_{1}$ versus $\Phi_{\mathrm{ext}}$ for qubit G, with a minimum frequency ($f_{01}$) of 197 MHz. Solid and dashed lines represent the dielectric loss model (Eq. 1) predictions at $T=0$ and $T=18.7$ mK, respectively. Inset: Qubit spectrum near flux frustration point. $\bf{b}$ Pure dephasing times $T_{\varphi}$ versus $\Phi_{\mathrm{ext}}$ for the same qubit, using free induction (Ramsey) and various dynamical decoupling sequences (spin echo, CPMG(2), CPMG(4), CPMG(8)). The black lines indicate dephasing times due to flux noise derived from a joint fit of all data. $\bf{c}$ $\zeta$ versus $f_{01}$ for our qubits in Table 1 (blue dots) and the $E_{J}$-tunable qubit from Ref. [31] (orange dots). Green and red solid lines represent $\zeta(f_{01})$ calculated using the dielectric loss model (Eq. 1) at $T=18.7$ mK and $T=0$, respectively. Predictions using the flux noise model with different $A_{\Phi}$ are shown as dashed lines. The grey shaded area signifies predicted $\zeta(f_{01})$ combining dielectric loss ($\tan\delta_{C}\in(1.21,4.93)\times 10^{-6}$ at $T=0$) and flux noise at $A_{\Phi}=2.43~{}\mu\Phi_{0}/\sqrt{\text{Hz}}$.
@@
In Fig. 4a and b, we present the measured $T_{1}$ and pure dephasing $T_{\varphi}$ times of qubit G as functions of $\Phi_{\text{ext}}$, respectively. Each data point for $T_{1}$ is extracted from fitting the relaxation data with an exponential decay curve at a specific flux position. The $T_{1}$ versus $\Phi_{\text{ext}}$ data exhibit a plateau near the flux frustration position and demonstrate suppression at $\Phi_{\text{ext}}=\Phi_{0}/2\pm 0.005\Phi_{0}$, likely due to interaction with a TLS at the corresponding qubit frequency. This flux dependence of $T_{1}$ precludes inductive-like loss channels [35]. This data is subsequently fitted using two variants of the dielectric loss model to determine the dielectric loss tangent, $\tan\delta_{C}$. One variant adheres to the widely used phenomenological dielectric loss model [2, 3]:
@@
dependent on the qubit’s finite effective temperature $T$. The alternative model posits a bath of TLS weakly interacting with the qubit, demonstrating temperature independent loss by taking the saturation effect of the TLS at finite temperature into account [31]. This model coincides with Eq. 1 at $T\rightarrow 0$, where the $\coth$ term converges to unity. The choice of model influences the derived loss tangent $\tan\delta_{C}$, typically yielding a higher value for the TLS bath model when the qubit energy $\hbar\omega_{01}\lesssim k_{B}T$. The qubit parameters such as the matrix element $\langle 0|\hat{\varphi}|1\rangle$, the charging energy $E_{C}$, and the qubit frequency $\omega_{01}=2\pi f_{01}$ are independently obtained from the qubit spectroscopy, where a representative spectrum of the corresponding qubit around the flux frustration position ($\Phi_{\text{ext}}=\Phi_{0}/2$) are shown in the inset of the Fig. 4a. We observe occasional suppression of $T_{1}$ at $\Phi_{\text{ext}}=\Phi_{0}/2$, suggestive of a TLS drifting into resonance with the qubit, when the qubit is idled at the same specific external flux position ($\Phi_{\text{idle}}=\Phi_{0}/2$). This suppression, however, can be mitigated by idling the qubit at $\Phi_{\text{idle}}\neq\Phi_{0}/2$ (see Appendix D for details). For instance, the data in Fig. 4a is collected with the qubit idled at $\Phi_{\text{idle}}=0.505\Phi_{0}$. This behavior might stem from alterations in the TLS environment triggered by qubit dynamics [36], presenting an intriguing avenue for future research.
In Fig. 4b, we present the typical qubit pure dephasing times $T_{\varphi}$ as a function of $\Phi_{\text{ext}}$, obtained from different dynamical decoupling schemes that utilize Ramsey, spin echo, CPMG($N$[37, 38] pulse sequences (with $N=2,4,8$ representing the number of $\pi$ pulses). Each data point is the result of fitting a Gaussian decay to the dephasing [39, 40, 41]. At the flux frustration point ($\Phi_{\text{ext}}=\Phi_{0}/2$), where the qubit exhibits first-order insensitivity to flux noise, we record the maximum $T_{\varphi}$. This dephasing time $T_{\varphi}$ decreases monotonically with deviation from this optimal position, a behavior typical of flux-type qubits [39]. The complete $T_{\varphi}$ versus $\Phi_{\mathrm{ext}}$ dataset, including Ramsey, spin echo, and CPMG($N$) sequences, is then fitted to a $1/f$ flux noise model [40], i.e. $S_{\Phi}(\omega)=A_{\Phi}^{2}(2\pi\times 1~{}\text{Hz}/\omega)$ for the extraction of the flux noise amplitude $A_{\Phi}$ (see Appendix F for more details).
@@
Table 1: Summary of device parameters and coherence metrics. $T_{1}$ and $T_{2,\mathrm{echo}}$ values, are obtained by fitting relaxation and dephasing to exponential decays at the flux frustration position. Asterisks (*) highlight observations of non-exponential $T_{1}$ decay. For qubit G, $\tan\delta_{C}$ and $A_{\Phi}$ are extracted from coherence time data as a function of $\Phi_{\text{ext}}$, collected with $\Phi_{\text{idle}}=0.505\Phi_{0}$. Coherence times versus $\Phi_{\text{ext}}$ for qubit H were not collected.